Session: May 2004 (session 2)
Wavelength: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Noto (P. Cassaro at Wed May 26 15:55:57 2004) Minor failures :
The whole observation has been carried out without problems, but the active surface has not been turned on. Possible bad data.

Onsala (Michael Lindqvist at Wed May 26 16:28:33 2004) Success :
We had some rain.

Jodrell_Bank (Alastair Gunn at Thu May 27 05:36:20 2004) Success :
The Lovell Telescope was used. No known problems. The schedule was manually changed to limit the number of calibrator scans on the Lovell telescope.

Jodrell_Bank (Alastair Gunn at Thu May 27 05:37:15 2004) Success :
MkII was used. No known problems.

Effelsberg (A. Kraus at Thu May 27 08:20:24 2004) Minor failures :
problems with the formatter. Data lost until UT0935.

Medicina (Andrea Orlati at Wed Jun 16 09:04:40 2004) Success :